(2t^2-5t-4)(3t^2-7+1/2) Multiply and show work. Check Answer.


This is what I did and the teacher says its still wrong....please help?
(2t^2-5t-4)
(3t^2-6.5)
6t^4-15t^3-12t-13t^2+32.5t+26
6t^4-15t^3-13t^2+20.5t+26

Answer by richwmiller(17219)

you could factor
1/2(2t^2-5t-4)*(6t^2-13)
but the final result is
6t^4-15t^3-25t^2+65t/2+26
here we have the various parts
(2t^2-5t-4)(3t^2)+(2t^2-5t-4)(-6.5)
2t^2(3t^2)=6t^4
-5t(3t^2)=-15t^3
-4(3t^2)=-12t^2
2t^2(-6.5)=-13t^2
add -25t^2
-5t(-6.5)+32.5t=65t/2
-4(-6.5)+26
your mistake was you had 12t and not 12t^2
6t^4-15t^3-12t-13t^2+32.5t+26